OpenStudy (anonymous):

Please show me working out for 4 1/4 - 1 3/9

OpenStudy (anonymous):

convert them to improper fractions first oh you can simplify the second one - it same as 1 1/3 4 1/4 = 17/4 and 1 1/3 = 4/3 17 - 4 -- -- 4 3 now you need to make the bottom parts (denominators) equal by finding the lcm (lowest common multiple) of 3 and 4 which is 12 so the sum becomes 51 - 16 35 -- -- = -- 12 12 12 answer is 35/12 or 2 11/12

OpenStudy (kira_yamato):

4 1/4 - 1 3/9 = 3 1/4 - 1/3 = 3 3/12 - 4/12 = 3 - 1/12 = 2 11/12